# Localized Distributional Robustness in Submodular Multi-Task Subset Selection

### v2

@@ 4 Relative Entropy Regularization
    for all $\lambda>0$ and $x\in[0,h(S^{\ast})]$, hence $g$ is a convex function.
@@
    As per (v), the first-order derivative of $g$ is a monotone increasing function over all of its domain, which attains its maximum value at $x=h(S^{\ast})$. Hence,
    $\dfrac{\mathrm{d}g}{\mathrm{d}x}\leq\dfrac{\lambda}{1-h(S^{\ast})},$ (46)
    making $g$ Lipschitz continuous with Lipschitz constant $\lambda/(1-h(S^{\ast}))$.
This theorem proves that standard methods such as Stochastic Greedy (presented in Algorithm 2), with well-established theoretical guarantees such as that of Theorem 1 are well-suited for the solution of Problem (25) (and hence for its equivalent, Problem 14). This is due to the fact that $g$ is a monotone increasing function, and that the maximizer of $g$ over its domain is $h(S^{\ast})$, as observed in the proof of the previous Theorem. Hence, any method approximating $h(S^{\ast})$ may be employed in maximizing $G$ as well. Regarding the approximation guarantee provided by Stochastic Greedy for the function $G(S)=g(h(S))$, we state the following Theorem.
@@
The set $\hat{S}$ constructed by Stochastic Greedy achieves, on the function $G$,
$\mathbb{E}[G(\hat{S})]\geq G(S^{\ast})-(1/e+\epsilon)\dfrac{h(S^{\ast})}{1-h(S^{\ast})}.$ (47)
@@
$\mathbb{E}[h(\hat{S})]\geq(1-1/e-\epsilon)h(S^{\ast}).$ (48)
@@
$\begin{split}h(S^{\ast})-\mathbb{E}[h(\hat{S})]&\leq h(S^{\ast})-(1-1/e-\epsilon)h(S^{\ast})=(1/e+\epsilon)h(S^{\ast}).\end{split}$ (49)
Now, by the Lipschitz continuity of $g$ with Lipschitz constant $\lambda/(1-h(S^{\ast}))$ (as per Theorem 2 , (vi)), we have
$g(h(S^{\ast}))-g(\mathbb{E}[h(\hat{S})])\leq(h(S^{\ast})-\mathbb{E}[h(\hat{S})])\dfrac{\lambda}{1-h(S^{\ast})}.$ (50)
@@
$g\left(h(S^{\ast})\right)-\mathbb{E}[g(h(\hat{S}))]\leq g(h(S^{\ast}))-g(\mathbb{E}[h(\hat{S})]).$ (51)
Then, combining (50) and (51), and rearranging, we have
$\mathbb{E}[G(\hat{S})]\geq G(S^{\ast})-(1/e+\epsilon)\dfrac{h(S^{\ast})}{1-h(S^{\ast})}.$ (52)
This, in turn, guarantees the production of a locally distributionally robust solution that is produced at no additional function evaluation cost with respect to the solution of the naive formulation of Problem (4). Our novel formulation greatly relaxes the SSA and enablexs the use of much less computationally expensive methods for the solution. Namely, Stochastic Greedy has a function evaluation cost of $\mathcal{O}(K\cdot\lvert R\rvert),$ compared to SSA’s $\mathcal{O}(\lvert N\rvert^{2}n\log(n\min_{i\in[n]}f^{i}(N)).$
Furthermore, this result extends gracefully to the case where the component objective functions $f^{i}$ for $i\in[n]$ are not submodular, but weak-submodular. Indeed, in this case, akin to the result with submodular functions, the newly-defined function $g$ in (25) turns out to be weak-submodular itself. The next Lemma demonstrates this result.
@@
$G(S)=-\lambda\log(\sum_{i=1}^{n}Q_{i}\exp{(\frac{-f^{i}(S)}{\lambda})})$ is weak-submodular when the component functions $f^{i}$ for $i\in[n]$ are weak-submodular.
@@
$\begin{split}w_{G}&=\max_{(S,T,e)\in\tilde{N}}\dfrac{\Delta_{G}(e|T)}{\Delta_{G}(e|S)}\\ &=\max_{(S,T,e)\in\tilde{N}}\dfrac{\log\left(\dfrac{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(T)}{\lambda}\right)}{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(T\cup\{e\})}{\lambda}\right)}\right)}{\log\left(\dfrac{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(S)}{\lambda}\right)}{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(S\cup\{e\})}{\lambda}\right)}\right)}.\end{split}$ (53)
@@
$\dfrac{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(S)}{\lambda}\right)}{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(S\cup\{e\})}{\lambda}\right)}=1,$ (54)
@@
$\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(S)}{\lambda}\right)=\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f^{j}(S\cup\{e\})}{\lambda}\right).$ (55)
@@
$\sum_{j=1}^{n}Q_{j}\underbrace{\left(\exp\left(\dfrac{-f^{j}(S)}{\lambda}\right)-\exp\left(\dfrac{-f^{j}(S\cup\{e\})}{\lambda}\right)\right)}_{A_{j}}=0.$ (56)
@@
Note that because each $f^{j}$ is monotone nondecreasing, $A_{j}$ in the sum is nonnegative for all $j$. Now, because $\sum_{j=1}^{n}Q_{j}=1$, there exists at least one $j$ such that $Q_{j}>0$. Let $j_{0}$ be one such $j$. We know that $A_{j_{0}}$ is either positive or zero. If it is the case that $A_{j_{0}}$ is positive, then there is at least one positive term in the sum (amongst other possibly positive, or zero terms), and hence the sum cannot be equal to zero. If, on the other hand, $A_{j_{0}}$ is zero, then we must have that
$f^{j_{0}}(S)=f^{j_{0}}(S\cup\{e\}).$ (57)
Yet, if this were the case, the WSC of $f^{j_{0}}$ would have been:
$\begin{split}w_{f^{j_{0}}}&=\max_{(S,T,e)\in\tilde{N}}\dfrac{\Delta_{f^{j_{0}}}(e|T)}{\Delta_{f^{j_{0}}}(e|S)}\\ &=\dfrac{\Delta_{f^{j_{0}}}(e|T)}{f^{j_{0}}(S\cup\{e\})-f^{j_{0}}(S)}\\ &=\dfrac{\Delta_{f^{j_{0}}}(e|T)}{0}=\infty,\end{split}$ (58)
@@
regardless of the choice of $T$ in the numerator. However, by the initial supposition that $f^{j}$ is weak-submodular for each $j\in[n]$, we know that $f^{j_{0}}<\infty$, and hence $A_{j_{0}}$ cannot be zero. Then, (56) cannot hold, and thus $w_{G}<\infty$, establishing that $G$ is a weak-submodular function. ∎
@@
This last result makes available to us theoretical guarantees presented in literature that demonstrate the efficacy of Stochastic Greedy on weak-submodular functions[39, 43].
@@
With all of these results settled, the only point that remains to be examined is whether solving the regularized, “soft-constrained” Problem (6), with which the entirety of this work has been interested, translates to solving the “hard-constrained” original Problem (5). We first note, as we previously have, the similarity of the formulation of Problem (6) to the Lagrangian of the inner minimization objective in Problem (5). Indeed, if one were to write the Lagrangian of the latter, one would have
~$\tilde{\mathcal{L}}(P,S,\lambda)=\sum^{n}_{i=1}P_{i}f^{i}(S)+\lambda D_{\text{KL}}(P\;\|\;Q)-\lambda R.$ (59)
@@
It is clear to see that the main difference between this Lagrangian and Problem (6) is the inclusion of the term $-\lambda R$. Now, in conventional fashion, in finding the dual of this problem, one would first minimize the Lagrangian with respect to the primal variable $P$, in terms of the dual variable $\lambda$, and obtain
$\tilde{G}(S,\lambda)\vcentcolon=\sum^{n}_{i=1}P^{\ast}_{i}(\lambda)f^{i}(S)+\lambda D_{\text{KL}}(P^{\ast}(\lambda)\;\|\;Q)-\lambda R.$ (60)
@@
Fortunately, the derivation starting in (17) remains valid, and so does the expression for $P^{\ast}$, given in (30), even with the addition of the $-\lambda R$ term. Then, continuing in the conventional fashion, one would maximize $\tilde{G}(S)$ with respect to $\lambda$ to obtain $\lambda^{\ast}$, which would, in turn, be used in the expression for $P^{\ast}$, to find the optimal value for the primal variable, $P^{\ast}(\lambda^{\ast})$, this time free of the variable $\lambda$.
@@
Let us for a moment adopt this approach in solving the inner minimization objective in Problem (6). In maximizing
@@
$\begin{split}\sum^{n}_{i=1}P^{\ast}_{i}(\lambda)f^{i}(S)+\lambda D_{\text{KL}}(P^{\ast}(\lambda)\;\|\;Q)=-\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}\right)\end{split}$ (61)
@@
with respect to $\lambda$, we swiftly realize that for $\lambda\geq 0$, the function is monotone increasing over all of its domain, and hence the maximizing $\lambda^{\ast}$ is unbounded. This, in turn, as the analysis in (28) shows, forces $P^{\ast}=Q$.
@@
This is precisely due to the missing $-\lambda R$ term. The inclusion of this term, in the “true Lagrangian” $\tilde{\mathcal{L}}$ given in (59), is the key to having a finite optimal $\lambda^{\ast}$, and hence having $P^{\ast}\neq Q$. The inclusion of this term makes both $\lambda^{\ast}$ and $P^{\ast}$ dependent on $R$, the radius of the neighborhood of robustness:
$\lambda^{\ast}=\arg\!\max_{\lambda\geq 0}-\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f^{i}(S)}{\lambda}\right)}\right)-\lambda R.$ (62)
@@
Because (62) is a concave maximization problem, taking a derivative of the objective function with respect to $\lambda$ and setting it equal to 00, we obtain the relation between $\lambda^{\ast}$ and $R$:
$\begin{split}&\dfrac{\sum_{i=1}^{n}Q_{i}\exp\left(\dfrac{-f^{i}(S)}{\lambda^{\ast}}\right)f^{i}(S)}{\lambda^{\ast}\sum_{i=1}^{n}Q_{i}\exp\left(\dfrac{-f^{i}(S)}{\lambda^{\ast}}\right)}+\log\left(\sum_{i=1}^{n}Q_{i}\exp\left(\dfrac{-f^{i}(S)}{\lambda^{\ast}}\right)\right)+R=0.\end{split}$ (63)
@@
With an appeal to the Implicit Function Theorem, we can say that within a neighborhood of any $(\lambda^{\ast}_{0},R_{0})$ pair that satisfies this relation, one can express $R$ as a one-to-one (hence invertible) function of $\lambda^{\ast}$, say, $R=r(\lambda^{\ast})$.
@@
The conclusion to these observations is the following: We claim that setting the regularization hyperparameter $\lambda$ to a value, say to $\lambda_{0}$, in the formulation of Problem (6) is akin to setting the $R$ in Problem (5) such that $R=r(\lambda_{0})$. In other words, one can envision changing the value of the regularization parameter $\lambda$ in the “soft-constrained” problem as equivalent to changing the value of the radius of the neighborhood of robustness in the “hard-constrained” problem. This is further supported by the fact that the $\lambda^{\ast}$ values obtained through the relation (63) as $R\to 0^{+}$ approach $\infty$, and those obtained as $R\to\infty$ approach 00.

### v3

@@ 4 Relative-Entropy Regularization
    for all $\lambda>0$ and $x\in[0,h(\mathcal{S}^{\ast})]$, hence $g$ is a convex function.
@@
    As per (v), the first-order derivative of $g$ is a monotone increasing function over its domain, attaining its maximum value at $x=h(\mathcal{S}^{\ast})$. Hence,
    $\left\lvert\dfrac{\mathrm{d}g}{\mathrm{d}x}\right\rvert\leq\dfrac{\lambda}{1-h(\mathcal{S}^{\ast})},$ (46)
    making $g$ Lipschitz continuous with constant $\lambda/(1-h(\mathcal{S}^{\ast}))$.
This theorem proves that standard methods such as Stochastic Greedy (presented in Algorithm 2), with well-established theoretical guarantees such as that of Theorem 1 are suitable for the solution of (25) (and hence for its equivalent, 14). This is due to the fact that $g$ is a monotone increasing function, and that the maximizer of $g$ over its domain is $h(\mathcal{S}^{\ast})$. Hence, any method approximating $h(\mathcal{S}^{\ast})$ may be employed in maximizing $G$ as well. Regarding the approximation guarantee provided by Stochastic Greedy for the function $G(\mathcal{S})=g(h(\mathcal{S}))$, we state the following theorem.
@@
The set $\mathcal{S}$ constructed by Stochastic Greedy in the solution of (25) satisfies
$\mathbb{E}[G(\mathcal{S})]\geq G(\mathcal{S}^{\ast})-(1/e+\epsilon)\dfrac{h(\mathcal{S}^{\ast})}{1-h(\mathcal{S}^{\ast})}.$ (47)
@@
$\mathbb{E}[h(\mathcal{S})]\geq(1-1/e-\epsilon)h(\mathcal{S}^{\ast}).$ (48)
@@
$\begin{split}h(\mathcal{S}^{\ast})-\mathbb{E}[h(\mathcal{S})]&\leq h(\mathcal{S}^{\ast})-(1-1/e-\epsilon)h(\mathcal{S}^{\ast})=(1/e+\epsilon)h(\mathcal{S}^{\ast}).\end{split}$ (49)
Now, by the Lipschitz continuity of $g$ with constant $\lambda/(1-h(\mathcal{S}^{\ast}))$ (as per Theorem 2, (vi)), we have
$g(h(\mathcal{S}^{\ast}))-g(\mathbb{E}[h(\mathcal{S})])\leq(h(\mathcal{S}^{\ast})-\mathbb{E}[h(\mathcal{S})])\dfrac{\lambda}{1-h(\mathcal{S}^{\ast})}.$ (50)
@@
$g\left(h(\mathcal{S}^{\ast})\right)-\mathbb{E}[g(h(\mathcal{S}))]\leq g(h(\mathcal{S}^{\ast}))-g(\mathbb{E}[h(\mathcal{S})]).$ (51)
Then, combining (50) and (51) and rearranging, we have
$\mathbb{E}[G(\mathcal{S})]\geq G(\mathcal{S}^{\ast})-(1/e+\epsilon)\dfrac{h(\mathcal{S}^{\ast})}{1-h(\mathcal{S}^{\ast})}.$ (52)
[…]
Furthermore, this result extends gracefully to the case where the component objective functions $f_{i}$ for $i\in[n]$ are not submodular, but weak submodular instead. In this case, akin to the result with submodular functions, the newly-defined function $g$ in (25) turns out to be weak submodular. The next lemma demonstrates this result.
@@
$G(\mathcal{S})=-\lambda\log(\sum_{i=1}^{n}Q_{i}\exp{(\frac{-f_{i}(\mathcal{S})}{\lambda})})$ is weak submodular when the component functions $f_{i}$ are weak submodular for all $i\in[n]$.
@@
$\begin{split}w_{G}&=\max_{(\mathcal{S},\mathcal{T},e)\in\tilde{\mathcal{N}}}\dfrac{\Delta_{G}(e\mid\mathcal{T})}{\Delta_{G}(e\mid\mathcal{S})}=\max_{(\mathcal{S},\mathcal{T},e)\in\tilde{\mathcal{N}}}\dfrac{\log\left(\dfrac{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{T})}{\lambda}\right)}{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{T}\cup\{e\})}{\lambda}\right)}\right)}{\log\left(\dfrac{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{S})}{\lambda}\right)}{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{S}\cup\{e\})}{\lambda}\right)}\right)}.\end{split}$ (53)
@@
$\dfrac{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{S})}{\lambda}\right)}{\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{S}\cup\{e\})}{\lambda}\right)}=1,$ (54)
@@
$\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{S})}{\lambda}\right)=\sum_{j=1}^{n}Q_{j}\exp\left(\dfrac{-f_{j}(\mathcal{S}\cup\{e\})}{\lambda}\right).$ (55)
@@
$\sum_{j=1}^{n}Q_{j}\underbrace{\left(\exp\left(\dfrac{-f_{j}(\mathcal{S})}{\lambda}\right)-\exp\left(\dfrac{-f_{j}(\mathcal{S}\cup\{e\})}{\lambda}\right)\right)}_{A_{j}}=0.$ (56)
@@
Note that because each $f_{j}$ is monotone nondecreasing, the $A_{j}$ term in the sum is nonnegative for all $j$. Now, because $\sum_{j=1}^{n}Q_{j}=1$, there exists at least one $j$ such that $Q_{j}>0$. Let $j_{0}$ be one such $j$. We know that $A_{j_{0}}$ is either positive or zero. If it is the case that $A_{j_{0}}$ is positive, then there is at least one positive term in the sum (alongside other nonnegative terms), and hence the sum cannot be equal to zero. If, on the other hand, $A_{j_{0}}$ is zero, then we must have that
$f_{j_{0}}(\mathcal{S})=f_{j_{0}}(\mathcal{S}\cup\{e\}).$ (57)
Yet, if this were the case, the WSC of $f_{j_{0}}$ would be
$\begin{split}w_{f_{j_{0}}}&=\max_{(\mathcal{S},\mathcal{T},e)\in\tilde{\mathcal{N}}}\dfrac{\Delta_{f_{j_{0}}}(e\mid\mathcal{T})}{\Delta_{f_{j_{0}}}(e\mid\mathcal{S})}=\dfrac{\Delta_{f_{j_{0}}}(e\mid\mathcal{T})}{f_{j_{0}}(S\cup\{e\})-f_{j_{0}}(\mathcal{S})}=\dfrac{\Delta_{f_{j_{0}}}(e\mid\mathcal{T})}{0}=\infty,\end{split}$ (58)
@@
regardless of the choice of $\mathcal{T}$ in the numerator. However, by the initial supposition that $f_{j}$ is weak-submodular for each $j\in[n]$, we know that $f_{j_{0}}<\infty$, and hence $A_{j_{0}}$ cannot be zero. Then, (56) cannot hold, and thus $w_{G}<\infty$, establishing that $G$ is a weak-submodular function. ∎
@@
This last result makes available to us theoretical guarantees related to Stochastic Greedy on weak submodular functions[25, 30].
@@
With all of this settled, one point that remains to be examined is whether solving the regularized, “soft-constrained” problem (6), with which the entirety of this work has been interested, translates to solving the “hard-constrained” original problem (5). We first note, as we previously have, the similarity of the formulation of (6) to the Lagrangian of the inner minimization objective in (5). Indeed, if we were to write the Lagrangian of the latter, we would have
$\tilde{\mathcal{L}}(P,\mathcal{S},\lambda)=\sum^{n}_{i=1}P_{i}f_{i}(\mathcal{S})+\lambda D_{\text{KL}}(P\;\|\;Q)-\lambda R.$ (59)
@@
Hence, the main difference between this Lagrangian and (6) is the term $\lambda R$. In conventional fashion, in finding the dual of this problem, we would first minimize the Lagrangian with respect to the primal variable $P$ in terms of the dual variable $\lambda$, and obtain
$\tilde{G}(\mathcal{S},\lambda)\vcentcolon=\sum^{n}_{i=1}P^{\ast}_{i}(\lambda)f_{i}(\mathcal{S})+\lambda D_{\text{KL}}(P^{\ast}(\lambda)\;\|\;Q)-\lambda R.$ (60)
@@
Fortunately, the derivation starting in (17) remains valid, and so does the expression for $P^{\ast}$ given in (30), even with the addition of the $\lambda R$ term. Then, continuing in the expected fashion, we would maximize $\tilde{G}(\mathcal{S})$ with respect to $\lambda$ to obtain $\lambda^{\ast}$, which we would use in the expression for $P^{\ast}$, to find the optimal value for the primal variable, $P^{\ast}(\lambda^{\ast})$, this time free of the variable $\lambda$.
@@
Let us for a moment adopt this approach in solving the inner minimization objective in (6). In maximizing
@@
$\begin{split}\sum^{n}_{i=1}P^{\ast}_{i}(\lambda)f_{i}(\mathcal{S})+\lambda D_{\text{KL}}(P^{\ast}(\lambda)\;\|\;Q)=-\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}\right)\end{split}$ (61)
@@
with respect to $\lambda$, we quickly realize that for $\lambda\geq 0$, the function is monotone increasing over its domain, and hence the maximizing $\lambda^{\ast}$ is unbounded. This, in turn, as the analysis in (28) shows, forces $P^{\ast}=Q$.
@@
This is precisely due to the missing $\lambda R$ term. The inclusion of this term in the true Lagrangian $\tilde{\mathcal{L}}$ in (59) is the key to having a finite optimal $\lambda^{\ast}$, and hence having $P^{\ast}\neq Q$. The inclusion of this term makes both $\lambda^{\ast}$ and $P^{\ast}$ dependent on $R$, the radius of the neighborhood of robustness:
$\lambda^{\ast}=\arg\!\max_{\lambda\geq 0}\;-\lambda\log\left(\sum_{i=1}^{n}Q_{i}\exp{\left(\frac{-f_{i}(\mathcal{S})}{\lambda}\right)}\right)-\lambda R.$ (62)
@@
Because (62) is the maximization of a concave function, taking the derivative of the function with respect to $\lambda$ and setting it equal to 0, we obtain the relation between $\lambda^{\ast}$ and $R$:
$\begin{split}&\dfrac{\sum_{i=1}^{n}Q_{i}\exp\left(\dfrac{-f_{i}(\mathcal{S})}{\lambda^{\ast}}\right)f_{i}(\mathcal{S})}{\lambda^{\ast}\sum_{i=1}^{n}Q_{i}\exp\left(\dfrac{-f_{i}(\mathcal{S})}{\lambda^{\ast}}\right)}+\log\left(\sum_{i=1}^{n}Q_{i}\exp\left(\dfrac{-f_{i}(\mathcal{S})}{\lambda^{\ast}}\right)\right)+R=0.\end{split}$ (63)
@@
With an appeal to the implicit function theorem, we can (informally) say that within a neighborhood of any $(\lambda^{\ast}_{0},R_{0})$ pair that satisfies this relation, one can express $R$ as a one-to-one (hence invertible) function of $\lambda^{\ast}$, say, $R=r(\lambda^{\ast})$.
@@
The conclusion to these observations is the following: We claim that setting the regularization hyperparameter $\lambda$ to a value, say to $\lambda_{0}$, in (6) is similar to setting the $R$ in (5) such that $R=r(\lambda_{0})$. In other words, one can envision changing the value of the regularization parameter $\lambda$ in the “soft-constrained” problem as equivalent to changing the value of the radius of the neighborhood of robustness in the “hard-constrained” problem. This is intuitively supported by the observation that the $\lambda^{\ast}$ values obtained through the relation (63) as $R\to 0^{+}$ approach $\infty$, and those obtained as $R\to\infty$ approach 0.